OpenStudy (anonymous):

Rewrite the expression as an equivalent expression that does not contain powers of trigonometric functions greater than 1. cos^3 x

OpenStudy (anonymous):



OpenStudy (anonymous):

@Libniz

OpenStudy (anonymous):

i dont understand step 3 and 4..

OpenStudy (anonymous):

cos(x)^3= since cos(x)^2= 1/2(cos(2x)+1) 1/2(cos(2x)+1) cos(x)

OpenStudy (anonymous):

From Mathematica:\[\cos ^3(x)=\frac{1}{4} (3 \cos (x)+\cos (3 x)) \]

OpenStudy (anonymous):

A. (3/4) cos x + (1/4) cos 3x + cos 2x B. (3/4) cos x + (1/4)cos 3x C.(3/4) cos x - (1/4) cos 3x D.(3/4) cos x - (1/4) cos 3x - cos 2x

OpenStudy (anonymous):

B:\[\frac{3 \text{ Cos}[x]}{4}+\frac{1}{4} \text{Cos}[3 x] \]

OpenStudy (anonymous):

@robtobey could u explain why?

OpenStudy (anonymous):

\[cos(x)=\frac{1}{2}(e^{it}+e^{-it})\] \[cos^3(x)=\frac{1}{2}(e^{it}+e^{-it})^3\] foil \[\frac{3 e^{-i t}}{8}+\frac{3 e^{i t}}{8}+\frac{1}{8} e^{-3 i t}+\frac{1}{8} e^{3 i t}\] \[\frac{1}{2}\left(\frac{3 e^{-i t}}{4}+\frac{3 e^{i t}}{4}+\frac{1}{4} e^{-3 i t}+\frac{1}{4} e^{3 i t}\right)\] so it is \[\frac{3}{4} cos[x]+\frac{1}{4} cos[x]\]

OpenStudy (anonymous):

ok thank u! @libniz

OpenStudy (anonymous):

ok let me do this now

OpenStudy (anonymous):